OpenStudy (anonymous):

HELP PLEASE What is the equation of the line that passes through the point (1, 6) and has a slope of 3? y = 3x + 5 y = 3x + 3 y = 3x - 17 y = 3x - 5

OpenStudy (ivettef365):

you need the equation y = mx + b you have m = slope = 3, all you need is to find the value of b, you have points (1,6) that represents (x,y)

OpenStudy (ivettef365):

With this information can you find value of "b"

OpenStudy (anonymous):

i dont know how to plug it in ): @ivettef365

OpenStudy (ivettef365):

ok, you have y = mx + b 6 = 3(1) + b solve for be

OpenStudy (ivettef365):

I substitute the m = 3 the x = 1 and y = 6 do you see it ??

OpenStudy (anonymous):

kind of im having a hard time ): @ivettef365

OpenStudy (anonymous):

wait it is B?

OpenStudy (ivettef365):

ok you have formula y = mx + b (x,y) (1,6) points given m = slope = 3 (given) so I substitute on the formula y = mx + b, the y for 6, the x for 1 and the m for 3 and you get the following; 6 = (3)(1) + b

OpenStudy (ivettef365):

now the way you solve this is 6 = 3 + b -3 -3 ---------- 3 = b

OpenStudy (ivettef365):

so now you know the value of b = 3 and you know that the slope = 3 so you have y = 3x + 3

OpenStudy (anonymous):

I am so confused @ivettef365

OpenStudy (ivettef365):

what is it that you don't understand ?

OpenStudy (anonymous):

oh wait no i get it now aha thank you (:

OpenStudy (ivettef365):

yw :)